OpenStudy (nicole143):

Find the first five terms of the sequence. @mathmale Hello, would you be willing to lend me hand in showing me how to complete this type of equation? I did a similar one but it didn't have the where a1 = 20 on the end.

OpenStudy (nicole143):

\[a_{n} = \frac{ 1 }{ 2 }a _{n} - 1 \] Where a1 = 20

OpenStudy (mathmale):

\[a_{n} = \frac{ 1 }{ 2 }a _{n} - 1\rightarrow a _{n=} \frac{ 1 }{ 2 }a _{n-1}\] is a bit clearer. Here's how you explain this formula: "The nth term of the sequence in question is equal to one half of the previous term, that is, one half of the (n-1)th term."

OpenStudy (mathmale):

So, Nicole, if a-sub-1 is 20, what is a-sub-2? Hint: Multiply 20 by (1/2).

OpenStudy (mathmale):

\[a _{1}=20 \rightarrow a _{2}=\frac{ 1 }{ 2 }(20)=?\]

OpenStudy (nicole143):

10 Sorry, I left there for a bit.. @mathmale

OpenStudy (mathmale):

So: a-1 is 20; a-2 is 10 (half of a-1) a-3 is ?

OpenStudy (nicole143):

Yes, thank you!

OpenStudy (nicole143):

Wait, what about to -1 ? @mathmale

OpenStudy (mathmale):

I'm sorry: my notation is not clear. I'll draw this sequence instead: